# Tightly Coupled SLAM with Imprecise Architectural Plans

### v2

@@ Abstract
Robots navigating indoor environments often have access to architectural plans, which can serve as prior knowledge to enhance their localization and mapping capabilities. While some SLAM algorithms leverage these plans for global localization in real-world environments, they typically overlook a critical challenge: the “as-planned” architectural designs frequently deviate from the “as-built” real-world environments. To address this gap, we present a novel algorithm that tightly couples LIDAR-based simultaneous localization and mapping with architectural plans under the presence of deviations. Our method utilizes a multi-layered semantic representation to not only localize the robot, but also to estimate global alignment and structural deviations between “as-planned” and “as-built” environments in real-time. To validate our approach, we performed experiments in simulated and real datasets demonstrating robustness to structural deviations up to cm and . On average, our method achieves less localization error than baselines in simulated environments, while in real environments, the “as-built” 3D maps show lower average alignment error.
Paper Video: https://www.youtube.com/watch?v=eEZPkcpjWlM
I Introduction
@@
Prior information from architectural plans can be used to enhance the localization and mapping accuracy of mobile robots. However, real-world buildings rarely match their plans perfectly due to construction tolerances and modifications. Incorporating such imprecise prior information into robot navigation pipelines can introduce systematic errors, potentially damaging the localization and mapping accuracy, instead of helping to improve it. To address this issue, it is critical to tightly couple the Simultaneous Localization and Mapping (SLAM) pipeline with the architectural plans. And for this, one needs a unified representation for both “as-planned” and “as-built” environments, that captures not only the geometry but also the architectural semantics. This is what will allow us to match structural elements between the “as-planned” and “as-built” instances and estimate the deviations between the two.
@@
Although there are various environment representation techniques such as occupancy grids [1], surface maps [2], octomap [3] etc, none of them explicitly models the semantic and hierarchical information of environment, which is needed to identify the deviated structural elements. Recent approaches such as 3D scene graphs [4, 5] or Situational Graphs (S-Graphs) [6, 7], represent a robot’s environment in a compact and hierarchical manner, encoding high-level semantic abstractions (for example, walls and rooms) and their relationships (e.g., a set of walls forms a room). Herein, S-Graphs extend 3D scene graphs by merging geometric models of the environment generated by SLAM approaches with 3D scene graphs into a multi-layered jointly optimizable factor graph. This representation, combined with the prior information extracted from architectural plans, can be used to provide fast and efficient localization.
@@
Coupling of SLAM factor graph with architectural plans to match structural elements (walls, rooms) between planned and built environments.
Globally localizing the robot in imprecise architectural plans.
@@ II Related Works
Most localization techniques using prior information from architectural plans assume that the environments are built precisely according to the plans. One of the most commonly used localization techniques in 2D metric prior maps is Monte Carlo Localization (MCL) [9], [10] but it is not scalable to large-scale complex environments. Boniardi et al [11] use a technique that scales to more complex environments by aligning a scan-based map with CAD-based floor plans. OGM2PGM [12] also scales to larger environments by converting the 2D floor plan to an occupancy grid map (OGM) and using a pose-graph map (PGM) to localize the robot. UKFL [13] further enhances the localization accuracy using an unscented Kalman filter to localize the robot in 3D metric meshes. Recent techniques such as [14] exploit neural networks to localize the robot using an implicit neural representation of the floor plans. All of the above mentioned techniques primarily rely on geometric information, not using any possible semantic information available in the architectural plans, limiting their ability to reason about the environment beyond geometric features. This geometric-only approach fails to leverage crucial semantic cues like room types, door locations, and functional spaces that could help disambiguate similar-looking areas and improve localization accuracy. In addition, inaccuracies or outdated information in the floor plan can significantly affect the performance of these methods.
@@
To address these limitations, semantic-based localization techniques, such as Mendez et al. [15] use semantic cues from architectural plans and sensor information to improve localization accuracy. Boniardi et al. [16] exploit the semantics of the room in architectural plans to do robot localization by matching the detected rooms from sensor data. Wang et al. [17] leverage prelabeled architectural features, such as wall intersections and corners, as landmarks in floor plans, and match them with detection from sensor data to jointly perform mapping and localization. Zimmerman et al. [18], [19] use high-level semantic information in floor plans, derived from object detection, along with geometric data from 2D LiDAR to perform long-term robot localization in floor plans. Huan et al. [20] convert architectural plans into semantically enriched point cloud maps, followed by a coarse-to-fine localization process using ICP. Gao et al. [21] used neural networks to detect vertical elements from floor plans to do LiDAR based localization. These methods are prone to inaccuracies due to misidentification and errors in the pose estimate of semantic elements. Moreover, they treat semantic elements in isolation, failing to leverage the rich contextual information embedded in their spatial and functional relationships.
Recent work such as Shaheer et al. [8] exploit the topological relationship between semantic elements to localize the robot with respect to architectural plans. However, all the above mentioned approaches assume no deviations between the architectural plans and the actual environment.
@@
Some recent works leverage imprecise floor plans for localization. Boniardi et al. [22] integrate localization techniques to take advantage of the information embedded in the CAD drawing, and the real-world observations acquired during navigation, which may not be reflected in the floor plan. Li et al. [23] presented a 2D LiDAR-based localization system in imprecise floor plans using stochastic gradient descent (SGD) with a scan matching algorithm. Chan et al. [24] presented a 2D LiDAR-based localization in floor plans that integrates SLAM with MCL. Blum et al. [25] use neural networks for feature segmentation and combine them with LiDAR data for localization in imprecise floor plans. While these works demonstrate robot localization in inaccurate floor plans, they lack the capability to identify element-wise deviations between ”as-planned” and ”as-built” environments.
Despite advances in robot localization using imprecise floor plans, no existing approach combines localization with the estimation of structural deviations between “as-planned” and “as-built” environments, limiting the ability to assess construction accuracy in real time. In this work, we address this limitation by presenting a unified framework that enables both global localization and deviation estimation.
@@ III System Architecture
III-A Factor Graph Modelling
Our algorithm models both SLAM and architectural plans as hierarchical factor graphs (see Fig. 3) composed of the following elements:
Architectural Graph (A-Graph). Three-layered hierarchical factor graph model of the geometry, semantics, and topology of an environment, generated from its architectural plan. It models the environment “as-planned” by the architect.
Situational Graph (S-Graph). Four-layered hierarchical optimizable factor graph built online from 3D LiDAR and odometry measurements [6], [7] which models the “as-built” environment. It also includes the keyframes in addition to the geometry, semantics, and topology of the environment.
Deviations Informed Situational Graph (diS-Graph). diS-Graph is the result of couppling both A-Graph and S-Graph.
These graphs are organized in the following layers:
Origins Layer. Both S-Graph and A-Graph are defined in their own reference frame. The reference frame of S-Graph and A-Graph are called and respectively. Both the origin frames are coupled via a Transformation factor called .
Keyframes Layer. Only present in S-Graphs, this layer contains as nodes the robot poses .
Walls Layer. In an A-Graph, this layer’s nodes encode two semantic entities, namely wall-surfaces and walls . Every wall-surface is defined by four coefficients representing its normal orientation and distance to origin [6]. We assume that each wall has two planar wall-surfaces with opposite orientations and the separation between them is equal to the width of the wall. In S-Graphs, this layer contains only the wall-surfaces extracted from 3D LiDAR scans. The keyframes that observe such wall-surfaces are linked to them through pose-plane constraints. This layer of both graphs is coupled via wall deviation factors in diS-Graph.
Rooms Layer. In an A-Graph, this layer also encodes two semantic entities, namely Rooms consisting of four wall-surfaces and Doorways . Two rooms constrain a doorway, and a room is constrained by four walls. In S-Graphs, this layer contains rooms comprising either four wall-surfaces or two wall-surfaces, and does not contain doorways. This layer of both graphs is coupled via room deviation factors in diS-Graph.
Floors Layer. In both A-Graph and S-Graph this layer consists of a floor center node represented as , constraining all rooms present at that particular floor level.
More details on the type of constraints between the different elements of the graphs can be found in [8].
@@ III-B The Algorithm
Our algorithm has two stages, as shown in Fig. 2. In Stage-A, the A-Graph is first matched and coupled with the S-Graph, and then jointly optimized. Stage-A runs until it finds the initial match between the two graphs, and provides the initial estimates of both the transformation between graphs and potential deviations between their elements. Afterwards, in Stage-B the joint graph is continuously refined through incremental matching and optimization as the robot navigates the environment, incorporating newly detected semantic entities. This tight coupling enables continuous refinement of both the robot’s localization and the detected structural deviations during navigation.
Both stages of our algorithm consist of multiple processes, described briefly below and detailed in their respective sections:
Graph Matching (Section IV). The Global Graph Matching in Stage-A provides the first unique match, when it exists, between the S-Graph and A-Graph at room and wall-surface levels, accounting for potential deviations. In Stage-B, the Local Graph Matching extends the previously matched elements in the diS-Graph with newly detected elements following an incremental approach.
Graph Coupling (Section. V). It is performed for the candidates matched by graph matching. In Stage-A, graph coupling registers the origins of two graphs along with the semantic coupling (wall-surfaces and rooms), and explicit mapping of the deviation factors. Stage-B only performs the semantic coupling that incorporates the newly observed entities with explicit deviation factors.
Graph Optimization (Section. VI). It optimizes the coupled graph. Stage-A involves two steps: alternating optimization and joint optimization. Stage-B only does joint optimization. Alternating optimization estimates two types of variables: 1. The ransformation between the origins of the S-Graph and the A-Graph yielding global localization, 2. The possible deviations between the matched graph entities. In joint optimization the coupled graphs are jointly optimized given the initial global transformation and deviation estimates.
@@ IV Graph Matching
Our graph matching extends the method presented in [8]. In [8], a top-down potential candidate search between the A-Graph and S-Graph is performed by leveraging their hierarchical structures. To assess the overall consistency of each generated candidate, two verification steps are applied iteratively. First, the consistency of the node type and graph structure is verified. Second, geometric consistency (i.e. norm) is maintained over a certain consistency threshold [26].
It is worth noting that symmetries may occur due to a lack of information as the robot has not visited the whole environment and thus the algorithm requires more information to provide a unique match, or when a large part of a building is symmetric and a unique match could never be found.
@@ IV-A Deviation-based Candidates Search
The presence of deviations generates geometric inconsistencies that affect the aforementioned candidates’ checks. Concretely, a deviation in the position of a wall-surface implies a slight deviation in the center of its parent room as well. To handle this, we propose a two-stage search algorithm where we first search for non-deviated wall-surfaces followed by the inclusion of those that are deviated.
Non-Deviated Elements Search To handle potential deviations, our method relaxes the matching criteria by decreasing the consistency thresholds at each level. First, we apply relaxed consistency thresholds for the generation of room-to-room and room-to-wall-surface match candidates, to account for the induced room-center inconsistencies. Then, to exclude deviated wall-surfaces, we increase the threshold for wall-surface-to-wall-surface candidates.
Deviated Elements Detection To identify the deviated wall-surfaces which were not matched in the first stage but are connected to already matched rooms, we decrease the consistency threshold at wall-surface-to-wall-surface level.
To further speed up the candidate search, we incorporate the following information: Orphan Wall-Surfaces: We utilize wall-surfaces in the S-Graph without a parent room for the assessment of the geometrical consistency of the final match candidates at the wall-surface level. Ground Orientation: We exploit the ground plane normal in the A-Graph and the S-Graph, only allowing candidates with z-axis rotations.
Finally, the geometric consistency score provides a quantification of the probability of the deviation for each room and wall-surface, which is further used in the Graph Coupling step (Section V).
@@ IV-B Incremental Matching
To enhance the efficiency of the Graph Matching algorithm in [8], we propose an incremental approach with two stages (associated with stages A and B of the systems architecture of Fig. 2), namely Global Graph Matching (Fig. 4 left) and Local Graph Matching (Fig. 4 right), each executing the two previously described deviation-based candidate search stages. Until a first unique match has been found, the Global Graph Matching is executed for every new observation in the S-Graph. Afterward, the Local Graph Matching is executed every time the diS-Graph is updated with newly observed rooms and wall surfaces. Here, already-matched elements are excluded from candidate generation, and each assessment of intra-level consistency considers the previously matched elements at the corresponding level.
@@ V Graph Coupling
Origin Coupling. We couple the origins of two graphs by introducing a transformation factor . The cost function is defined as:
Here and are the origins of the A-Graph and S-Graph respectively, and is the transformation between them. stands for the covariance of the cost, and it is always assigned a high value to estimate the transformation factor accurately.
Semantic Coupling. We couple the wall-surfaces and rooms of the A-Graph and the S-Graph to estimate deviations between the two graphs by introducing deviation factors as follows:
Room Coupling: To estimate the deviation in the pose of a matched room between the two graphs, we define a deviation factor between the two rooms as , where the cost function is defined as:
Here and are rooms of A-Graph and S-Graph, each consisting of a set of four wall-surfaces. is the covariance associated with the cost function depending on the probability of deviation estimated by the matching of the graph (Section IV). Rooms with a higher probability of deviation assigned by graph matching have a higher covariance assigned to their cost function than rooms with lower deviation probability. If all matched rooms have the same deviation probability, they are assigned lower uniform covariances.
Wall-Surface Coupling: After coupling the rooms of the A-Graph and the S-Graph, we then couple the wall-surfaces of the coupled rooms. We define the deviation factor between two wall surfaces as . The cost function to estimate the deviation value is defined as:
Here , where and are the normal orientation and distance of a plane in the A-Graph. Similarly, is a plane of the S-Graph with respect to the S-Graph origin. is the covariance associated with the cost function. Like rooms, wall-surfaces with a higher deviation are assigned higher covariances, and the ones with a lower deviation are assigned lower uniform covariances.

### v3

@@ Abstract
Robots navigating indoor environments often have access to architectural plans, which can serve as prior knowledge to enhance their localization and mapping capabilities. While some SLAM algorithms leverage these plans for global localization in real-world environments, they typically overlook a critical challenge: the “as-planned” architectural designs frequently deviate from the “as-built” real-world environments. To address this gap, we present a novel algorithm that tightly couples LIDAR-based simultaneous localization and mapping with architectural plans in the presence of deviations. Our method utilizes a multi-layered semantic representation to not only localize the robot, but also to estimate global alignment and structural deviations between “as-planned” and “as-built” environments in real-time. To validate our approach, we performed experiments in simulated and real datasets demonstrating robustness to structural deviations up to cm and . On average, our method achieves less localization error than baselines in simulated environments, while in real environments, the “as-built” 3D maps show lower average alignment error.
Paper Video: https://www.youtube.com/watch?v=9O0qwNhTuqk
Index Terms:
SLAM, Localization, Robotics and Automation in ConstructionI Introduction
@@
Prior information from architectural plans can be used to enhance the localization and mapping accuracy of mobile robots. However, real-world buildings rarely match their plans perfectly due to construction tolerances and modifications [1, 2]. Incorporating such imprecise prior information into robot navigation pipelines can introduce systematic errors, potentially damaging the localization and mapping accuracy, instead of helping to improve it. To address this issue, it is necessary to tightly couple the Simultaneous Localization and Mapping (SLAM) pipeline with architectural plans. This requires a unified representation of both the “as-planned” and “as-built” environments –one that captures not only the geometry but also architectural semantics. Such a representation will first enable the matching of structural elements between the two instances, and subsequently, the estimation of their deviations.
@@
Although there are various environment representation techniques such as occupancy grids [3], none of them explicitly models the semantic and hierarchical information of the environment, which is needed to identify the deviated structural elements. Recent approaches such as 3D scene graphs [4, 5] or Situational Graphs (S-Graphs) [6, 7], represent a robot’s environment in a compact and hierarchical manner, encoding high-level semantic abstractions (for example, walls and rooms) and their relationships (e.g., a set of walls forms a room). Herein, S-Graphs extend 3D scene graphs by merging geometric models of the environment generated by SLAM approaches with 3D scene graphs into a multi-layered jointly optimizable factor graph. This representation, combined with the prior information extracted from architectural plans, can be used to provide fast and efficient localization.
@@
Tightly coupling the SLAM factor graph with architectural plans, in order to match structural elements (walls, rooms) between planned and built environments.
Globally localizing the robot in imprecise architectural plans that may contain deviations from the real building.
@@ II Related Works
Most localization techniques using prior information from architectural plans assume that the environments are built precisely according to the plans. One of the most commonly used localization techniques in 2D metric prior maps is Monte Carlo Localization (MCL) [9], [10] but it is not scalable to large-scale complex environments. Boniardi et al [11] use a technique that scales to more complex environments by aligning a scan-based map with CAD-based floor plans. OGM2PGM [12] also scales to larger environments by converting the 2D floor plan to an occupancy grid map (OGM) and using a pose-graph map (PGM) to localize the robot. UKFL [13] further enhances the localization accuracy using an unscented Kalman filter to localize the robot in 3D metric meshes. Recent techniques such as [14] exploit neural networks to localize the robot using an implicit neural representation of the floor plans. These techniques rely solely on geometric information from architectural plans, ignoring available semantic data (room types, door locations, functional spaces) that could disambiguate similar areas and improve localization. Additionally, these geometric-only approaches are vulnerable to floor plan inaccuracies or outdated information.
@@
To address these limitations, semantic-based localization techniques, such as Mendez et al. [15] use semantic cues from architectural plans and sensor information to improve localization accuracy. Boniardi et al. [16] exploit the semantics of the room in architectural plans to do robot localization by matching the detected rooms from sensor data. Wang et al. [17] leverage prelabeled architectural features, such as wall intersections and corners, as landmarks in floor plans, and match them with detection from sensor data to jointly perform mapping and localization. Zimmerman et al. [18], [19] use high-level semantic information in floor plans, derived from object detection, along with geometric data from 2D LiDAR to perform long-term robot localization in floor plans. Huan et al. [20] convert architectural plans into semantically enriched point cloud maps, followed by a coarse-to-fine localization process using ICP. Gao et al. [21] used neural networks to detect vertical elements from floor plans to do LiDAR based localization. These methods are prone to inaccuracies due to misidentification and errors in the pose estimate of semantic elements. Moreover, they treat semantic elements in isolation, failing to leverage the rich contextual information embedded in their spatial and functional relationships. Recent work, such as Shaheer et al. [8] exploits the topological relationship between semantic elements to localize the robot with respect to architectural plans. However, all the above mentioned approaches assume no deviations between the architectural plans and the actual environment.
@@
Some recent works leverage imprecise floor plans for localization. Boniardi et al. [22] integrate localization techniques to take advantage of the information embedded in the CAD drawing, and the real-world observations acquired during navigation, which may not be reflected in the floor plan. Li et al. [23] presented a 2D LiDAR-based localization system in imprecise floor plans using stochastic gradient descent (SGD) with a scan matching algorithm. Chan et al. [24] presented a 2D LiDAR-based localization in floor plans that integrates SLAM with MCL. Blum et al. [25] use neural networks for feature segmentation and combine them with LiDAR data for localization in imprecise floor plans. Despite advances in robot localization using imprecise floor plans, no existing approach combines localization with the estimation of structural deviations between “as-planned” and “as-built” environments, limiting the ability to assess construction accuracy in real time. In this work, we address this limitation by presenting a unified framework that enables both global localization and deviation estimation.
@@ III System Architecture
III-A Factor Graph Structure and Notation
Our algorithm represents the SLAM backend and the architectural plans using hierarchical factor graphs (see Fig. 3). For clarity, we first introduce here the terminology and notation used throughout this paper.
Architectural Graph (A-Graph). Three-layered hierarchical factor graph containing the geometry, semantics, and topology of an environment, generated from its architectural plan [8]. It models the environment “as-planned” by the architect.
Situational Graph (S-Graph). Four-layered hierarchical optimizable factor graph, estimated online from 3D LiDAR and odometry measurements, that models the “as-built” environment [6], [7]. It also includes keyframes in addition to the geometry, semantics, and topology of the environment.
Deviations Informed Situational Graph (diS-Graph). Joint graph, coupling both the A-Graph and the S-Graph of a building. The layers of these graphs are as follows.
Origins Layer.
The S-Graph and A-Graph are defined in their respective frames and , with anchor frame set such that .
Keyframes Layer.
Exclusive to S-Graphs, this layer contains robot poses in the frame: .
Walls Layer. In A-Graph: Contains wall-surfaces and walls . Wall-surfaces are defined by their normals and distances to the origin [6]. A wall is constrained by two opposite wall-surfaces via the factor , defined in Eq. 2 of [8]. In S-Graph: Contains only wall-surfaces extracted from 3D LiDAR scans and linked to keyframes via pose-plane constraints. We couple both graphs in our diS-Graphs via wall deviation nodes and their associated factors , that we show in Fig. 3 and define in Eq. 3 in this paper.
Rooms Layer. In A-Graph: Contains rooms composed of four wall-surfaces, and doorways . A doorway constrains two rooms via the factor , that is defined in Eq. 3 of [8]. A room is constrained by four wall-surfaces via factor , defined in Eq. 7 of [7]. In S-Graph: Contains rooms of four wall-surfaces or two wall-surfaces . Both graphs are connected by room deviation factors in our diS-Graph.
Floors Layer. In both A-Graph and S-Graph this layer consists of a floor center node represented as and respectively, constraining all rooms present at that particular floor level with the factor , defined in Eq. 9 of the reference [7].
@@ III-B The Algorithm
Our algorithm has two stages, as shown in Fig. 2. In Stage-A, the A-Graph is first matched and coupled with the S-Graph, and then jointly optimized. Stage-A runs until it finds the initial match between the two graphs, and provides the initial estimates of both the transformation between graphs and potential deviations between their elements. Afterwards, in Stage-B, the joint graph is continuously refined through incremental matching and optimization as the robot navigates the environment, incorporating newly detected semantic entities. This tight coupling enables continuous refinement of both the robot’s localization and the detected structural deviations during navigation. Both stages of our algorithm consist of multiple processes, introduced briefly here and fully detailed later in their respective sections.
Graph Matching (Section IV). The Global Graph Matching in Stage-A provides the first unique match, when there is sufficient information, between the S-Graph and A-Graph at room and wall-surface levels, accounting for potential deviations. In Stage-B, the Local Graph Matching extends the previously matched elements in the diS-Graph with newly detected elements following an incremental approach.
Graph Coupling (Section. V). It is performed for the candidates matched by graph matching. In Stage-A, graph coupling registers the origins of two graphs along with the semantic coupling (wall-surfaces and rooms), and explicit mapping of the deviation factors. Stage-B only performs the semantic coupling between newly observed entities and their correspondences in the plan, with explicit deviation factors.
Graph Optimization (Section. VI). It optimizes the coupled diS-Graph. Stage-A involves two steps: alternating optimization and joint optimization. Alternating optimization estimates two types of variables, 1) The transformation between the origins of the S-Graph and the A-Graph, yielding global localization, and 2) the potential deviations between the matched graph entities. Stage-B jointly optimizes the diS-Graph from initial seeds for the global transformation and map deviations.
@@ IV Graph Matching
Our graph matching extends the method presented in [8]. In [8], a top-down potential candidate search between an A-Graph and a S-Graph is performed by leveraging their hierarchical structures. To assess the overall consistency of matching candidate, two verification steps are applied iteratively. First, the consistency of the node type and graph structure is verified. And secondly, geometric consistency is checked over a certain predefined threshold [26]. Finally, the consistencies of all candidates are compared, and the candidate with the best consistency is selected as the final unique match. Note that symmetries may affect our algorithm. Their occurrence may result in false positives when the robot has only partially explored the environment and observed one of the symmetric places. They may also cause several equally good candidates if the robot has explored a larger fraction of the building.
@@ IV-A Deviation-based Candidates Search
The presence of deviations generates geometric inconsistencies that affect the aforementioned candidates’ checks of [8]. Concretely, a deviation in the position of a wall-surface implies a slight deviation in the center of its parent room as well. To handle this, we propose a two-stage search method in this paper, searching first for non-deviated wall-surfaces and including after that those elements that are deviated.
Non-deviated wall-surfaces search To handle potential deviations, we tighten the matching criteria. Specifically, during the generation of non-deviated room-to-room and room-to-wall-surface match candidates, we look for the candidates with higher consistency at each level.
Deviated wall-surfaces detection To identify deviated wall-surfaces that were not matched in the first stage but are connected to already. matched rooms, we now realx the matching criteria at wall-surface-to-wall-surface level to account for inconsistencies due to deviations. To further speed up the candidate search, we incorporate the following additional information.
Orphan wall-surfaces: We utilize wall-surfaces in the S-Graph without a parent room for the assessment of the geometrical consistency of the final match candidates at the wall-surface level.
Ground orientation: We exploit the ground plane normal in the A-Graph and the S-Graph, only allowing candidates with z-axis rotations. Finally, the geometric consistency score provides a quantification of the probability of the deviation for each room and wall-surface, which is further used in the graph coupling stage (Section V).
@@ IV-B Incremental Matching
To enhance the efficiency of the Graph Matching algorithm in [8], we implement an incremental approach with two stages (associated with stages A and B of the systems architecture of Fig. 2), namely Global Graph Matching (Fig. 4 left) and Local Graph Matching (Fig. 4 right), each of them based on the two deviation-based candidate search stages described in Section IV-A. Until a first unique match has been found, the Global Graph Matching is executed for every new observation in the S-Graph. Afterwards, the Local Graph Matching is run every time the diS-Graph is updated with newly observed rooms and wall-surfaces. The graph elements that have been already matched are excluded from candidate generation here, and each assessment of intra-level consistency considers the previously matched elements at the corresponding level.
@@ V Graph Coupling
Origin Coupling. We couple the two graphs using the anchor node as a common reference frame, as shown in Fig. 3. This allows us to align both graphs in the frame by estimating the transformation between and . The associated constraint is defined as:
stands for the covariance of the cost, and it is assigned a high value to estimate the transformation accurately. Note that, since our approach simultaneously estimates both the global transformation and structural deviations, special care should be put in choosing the initial seed for this transformation and avoiding local minima coming from the combined effect of wall deviations and local registration.
Semantic Entities Coupling. To estimate deviations between the wall-surfaces and rooms of the A-Graph and the S-Graph, we couple them with deviation nodes and respectively, as shown in Fig. 3.
Room Coupling: To estimate the deviation in the pose of a pair of matched rooms of the two graphs, we define the cost function as:
Here and are the room poses of A-Graph and S-Graph respectively, in common frame . is the covariance associated with the cost function depending on the probability of deviation estimated by the matching of the graph (Section IV). Rooms with a higher probability of deviation assigned by graph matching have a higher covariance assigned to their cost function than rooms with lower deviation probability. If all matched rooms have the same deviation probability, they are assigned lower uniform covariances.
Wall-Surface Coupling: To estimate the deviation between a pair of matched wall-surfaces of both graphs, we define the cost function of deviation factor between two wall-surfaces as:
Here and are the poses of wall-surfaces of A-Graph and S-Graph respectively, in common frame .111For compactness, we slightly abused notation by overloading the and operators. In Eq. 2, they denote the pose composition and pose difference, respectively, and, in Eq. 3, they denote the plane compositions and difference, respectively. is the covariance associated with the cost function. Like rooms, wall-surfaces with a higher deviation are assigned higher covariances, and the ones with a lower deviation are assigned lower uniform covariances.